Write a balanced equation for the reaction between magnesium metal with hydrochloric acid. If you mix 6.51 g Mg with 45.6 mL of 6.00 M HCl, how many liters of hydrogen gas will you get at 25°C and 101 kPa? (limiting reactant problem:)) *
1
Expert's answer
2020-05-25T14:12:58-0400

Mg+2HCl=MgCl2+H2.

6.51/24 = 0,27125.

6 x 0.0456 /2= 0,1368.

V= 0.1368 x 8.314 x 298 / 101= 3,35576089 L.